OpenStudy (anonymous):

derivative of 4x^2 + 3y^2 + 16x + 36y - 68=0

OpenStudy (anonymous):

8x+6ydy/dx+16+36dy/dx=0

OpenStudy (anonymous):

from here finf the value of dy/dx..can u do it?

OpenStudy (anonymous):

-8x - 16 / 42

OpenStudy (08surya):

8x+6ydy/dx+16+36dy/dx=o dy/dx(6y+36)=-8x-16 dy/dx=(-8x-16)/(6y+36)

OpenStudy (anonymous):

thanksssssss

OpenStudy (anonymous):

-8x - 16 / 42